OpenStudy (anonymous):

David kicks a soccer ball off the ground and in the air, with an initial velocity of 35 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? 17.9 feet 18.2 feet 18.7 feet 19.1 feet @amistre64 @triciaal @confluxepic

OpenStudy (welshfella):

i dont think there's enough information given to answer this question

OpenStudy (anonymous):

well thats all the info i got.

OpenStudy (welshfella):

we dont know the time t , we know v and as the ball is on the ground we can take s to be 0

OpenStudy (anonymous):

okay.

OpenStudy (welshfella):

so we have h(t) = -16t^2 + 35t - 2 unknowns

OpenStudy (anonymous):

okay.

OpenStudy (welshfella):

oh if we differentiate we get d h / dt = -32t + 35 and this = 0 for maximum h so t = 35/32 seconds and this gives h as - 16 * (35/32)^2 + 35 * 35/32

OpenStudy (anonymous):

okay but im confused what do i do?

OpenStudy (welshfella):

hold on - have you done any calculus Finding derivatives?

OpenStudy (anonymous):

no

OpenStudy (welshfella):

oh ok

OpenStudy (welshfella):

well you can't do it that way then

OpenStudy (anonymous):

ohh

OpenStudy (welshfella):

we need to find the vertex of the graph -=16t^2 + 35t the graph looks something like this

OpenStudy (anonymous):

i can graph it

OpenStudy (welshfella):